The length of the top of a table is 5 ft greater than the width. Find the length and width if the area is 84 ft^2.
Answer by checkley77(12844)

L=W+5
LW=84
W(W+5)=84
W^2+5W-84=0
(W-7)(W+12)=0
W-7=0
W=7 ANS. FOR THE WIDTH.
L=7+5
12 ANS. FOR THE LENGTH.
PROOF:
7*12=84
84=84